Compute the interest rate at which $500 per month payments should grow to accumulate savings of $50,000 in 3 years. Payments are made at the end of each month. (Note: Round your answer to two decimal places.)

EFFECT

DISC

RATE

NOMINAL

A

B

1 Data Description
2 NPER 3 x 12 = 36
3 PMT 500
4 PV 0
5 FV 50,000
6 [Type] 0
7 Interest Rate =   
8 = %

Compute the stated interest rate on a promotional credit card if the effective interest rate is 8.0%. Assume that interest is compounded (a) semi-annually and (b) quarterly. (Note: Round your answers to two decimal places.)

NOMINAL

RATE

EFFECT

DISC

A

B

1 Data Description
2 Compounded semi-annually
3 Interest Rate =   
4 = %
5 Compounded quarterly
6 Interest Rate =   
7 = %
8

Compute the actual interest rate paid or earned on a mortgage loan if the bank states that the nominal rate is 10%, and the interest is compounded (a) semi-annually and (b) quarterly. (Note: Round your answers to two decimal places.)

EFFECT

DISC

RATE

NOMINAL

A

B

1 Data Description
2 Compounded semi-annually
3 Interest Rate =   
4 = %
5 Compounded quarterly
6 Interest Rate =   
7 = %
8
Answer #1

1.
=RATE(12*3,-500,0,50000,0)*12=62.37328%

2.
Semi-annually=(1+8%/2)^2-1=8.16%

Quarterly=(1+8%/4)^4-1=8.24%

3.
Semi-annually=(1+10%/2)^2-1=10.25%

Quarterly=(1+10%/4)^4-1=10.38%